OpenStudy (anonymous):

2x - 5y = -1 x - 3y = -3 I need to find the solution by using subsitution method

OpenStudy (phi):

As a first step, I would "solve for x" in the 2nd equation x - 3y = -3 can you do that?

OpenStudy (anonymous):

no, don't know where to begin

OpenStudy (anonymous):

how do I subsitute for x when I am not sure what I am subsituting

OpenStudy (phi):

When you get a chance http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/v/solving-for-a-variable There are more that show how to solve equations.

OpenStudy (phi):

But for this problem x - 3y = -3 add +3y to both sides of the equation x -3y +3y = -3 + 3y

OpenStudy (phi):

you can simplify the left side: if you have 3 y's and take away 3 y's you are left with no y's you get x= -3 +3y that means that you can replace x in the 1st equation with the equivalent (-3+3y) can you do that? what do you get?

OpenStudy (anonymous):

yes I can. I get 2(-3 +3y) - 5y = -1 -6 + 6y -5y = -1 -6 - y = -1 -6 + 6 -y = -1 + (-6) -y= -5

OpenStudy (anonymous):

I got x = -3 and y = -1

OpenStudy (phi):

you have the right idea, but need to be more careful. I get 2(-3 +3y) - 5y = -1 -6 + 6y -5y = -1 first notice you have 6 y take away 5 y how many y's do you get?

OpenStudy (anonymous):

I only have one y left

OpenStudy (anonymous):

for my answer I get x as -3 and y = -1

OpenStudy (phi):

yes so -6 + 6y -5y = -1 is -6 +y = -1 now if you add +6 to both sides, what do you get? +6-6+y= +6-1

OpenStudy (anonymous):

y = 5

OpenStudy (phi):

now use x= -3 +3y replace y with 5 and then figure out what x is

OpenStudy (anonymous):

x = -3 + 3(5) x = -3 + 15 x = 12

OpenStudy (phi):

ok now test if your answer works x=12, y =5 Does it work in the 1st equation? 2x - 5y = -1

OpenStudy (anonymous):

2(12) -5(5) = -1 24 - 25 =-1 -1 = -1 yes I understand thank you so much this made it easier for me to work my other problems. Your a life saver

OpenStudy (phi):

ok. The videos are helpful (for me anyways) if you need "inspiration"

OpenStudy (anonymous):

great thanks